calculate the area of a circle whose radius is 2cm
Answer by jim_thompson5910(35256)


A = pi*r^2

A = (3.14159)*(2)^2

A = (3.14159)*(4)

A = 12.56636

So the area is approximately 12.56636 square cm

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim